General instructions:

✓ Discuss the following points as much as possible.

✓ Each questions should be answered.

✓ Attempt all questions, size of the work 5-10 pages.

Submitted date:- Tus, Aug 15

Questions

 1. Show by using graph the factors which can able to affect the equilibrium conditions and explain how the government can able to inverse in the market?

2. In the market the inverse demand function given as P=200-1/6Qd and P=100+1/2Qs

If in the market there are 200 identified by years and 100 supplies based on the information: 

A. Calculate Pe and Qe

B. Calculate Ped and Pes 

C. When price increases in the market what will happen for TR

D. When the price value fixed in the market at $16

√ what will happen

√ whose problem do it

√ how it will be managed

√ number of demand and supply

3. Explain by using the graph the relation between TPI, API, MPI and stages of SR production?

4. Show price equilibrium and quantity of the demand and supply function?


1
Expert's answer
2021-08-09T06:49:55-0400

1) Change in supply and demand affect equilibrium conditions.

For example decrease in demand makes both price and quantity to fall while decrease in supply raises prices and lower quantity as shown below.



Government can control such situation by having price controls.


2)

"We=Qs\\200-\frac{1}{6}Q=100-\frac{1}{2}\\200-100=\frac{1}{6}+\frac{1}{2}\\Q=150\\p=100-\frac{1}{2}(150)=175"

Total revenue will increase when price rises.


3)


4)y=Mx+b

y=y intercept=price

b=x intercept=quantity